a disc is rotating with an angular velocity . a constant retarding torque is applied on it to stop the disc. the angular velocity becomes half after n rotaions. how many more rotations will it make before coming to rest? 

Grade:12

1 Answers

Akash Kumar Dutta
11 years ago

n/3 rotations,
apply w(f)^2 = w(i)^2 + 2.a.(2n.pi).....where a =acc and constant for the system.